For the function below, find a) the critical numbers; b) the open intervals where the function is increasing; and c) the openFor the function below, find a) the critical numbers; b) the open intervals where the function is increasing; and c) the openDetermine the location of each local extremum of the function. f(x) = -x2 - 9x2 - 15x+1 What is/are the local minimum/minima?
Answer #1

-0 ① ffa) = 2x +32_12x+2 critical points are given by f(x)=0 6 xt 62-12 zo azx-2=0 2²722-2x-2=0 x (2+2) -81 (2+2) =0 (2-1) (x3 -x q²_15x+1 f(x) f(x) zo ista)= -37-187-15 = - the critical points are sy - 322_188-1520 a +62+5 =0 =*+5x+x+520 à (4+5)+1(a